OpenStudy (anonymous):

Input in standard form the equation of the given line. The line with m = -3/2 and b = 6

OpenStudy (anonymous):

start with slope-intercept form: \[y=mx+b=(-3/2)x+6\]Standard form is:\[y+(3/2)x=6\]But we don't want the fraction so we multiply through by the denominator:\[3x+2y=12\]

OpenStudy (saifoo.khan):

y = -3x/2 + 6 LCM y = -3x + 12 ---------- 2 2y = -3x + 12